# Optical probes of two-component pairing states in transition metal dichalcogenides

### v1

@@ Abstract
$\sigma_{ij}=\frac{\pi e^{2}}{2\omega V}\sum_{m\neq n}v^{i}_{nm}v^{j}_{mn}f_{nm}\delta(\varepsilon_{n}-\varepsilon_{m}+\omega),$ (8)
@@
where $e$ is the electron charge, $V$ is the volume of the system, $f_{nm}=f(\varepsilon_{n})-f(\varepsilon_{m})$ is the difference of Fermi functions, which depend on the energy $\varepsilon_{n}=E_{n}-\mu$ of band $n$, the chemical potential $\mu$ and the inverse temperature $\beta=1/k_{B}T$ measured in units of the Boltzmann constant $k_{B}$. We set $\beta=300\mathrm{eV}^{-1}$ ($T\simeq 3.87$K) [^1] It is important to note that the calculation of the optical conductivity has no information about the transition temperature of the superconductor and it does not require to be under the critical temperature. Therefore the temperature used here just plays the role of a small softening of the Fermi functions.. The velocity matrices $v^{i}_{mn}$ account for the transition amplitude between states $n,m$ and can be defined as:
@@
Note the optical conductivity of clean multiband superconductorsAhn and Nagaosa (2021a) as computed here can be reduced in size for some types of superconductors -but not suppressed- by inclusion of vertex correctionsWatanabe and Watanabe (2024), which are outside of the scope of this work.
@@
Anisotropy in the nematic phase –– In the presence of $C_{3}$ symmetry, the optical conductivity satifies $\sigma_{xx}=\sigma_{yy}$. Since the two-component nematic pairing breaks this symmetry, we consider the difference $\sigma_{xx}-\sigma_{yy}$ as a probe of the nematic pairing. Fig. 3 shows this difference in the range of frequencies near the superconducting gap for both $\Delta_{p_{x}}$ and $\Delta_{p_{y}}$. A sizable conductivity occurs due to transitions near the Fermi level, with significant differences between $\Delta_{p_{x}}$ and $\Delta_{p_{y}}$ which can help determine the pairing anisotropy.
@@
As a further optical signature of the pairing in multiorbital systems, we also consider higher frequency effects which occur due to transitions from the Fermi level BdG quasiparticles to unoccupied electron states, and from occupied electron states to BdG quasiparticles. These occur in the frequency range where the optical conductivity of the normal state is finite (see Fig. 1 (c)) and represent transfer of spectral weight due to superconductivity as discussed in Ref. [Ahn and Nagaosa, 2021b]. However, here we consider responses which are zero in the absence of pairing, due to $C_{3}$ and time-reversal symmetry. The observation of any optical conductivity is therefore a signature of multiorbital pairing, which is shown in the inset of Fig. 3. The anisotropy of the nematic pairing is reflected in this spectrum as well.
@@
Figure 3: Anisotropy of the optical conductivity at low (main figures) and high (inset) frequencies for the nematic phases $\Delta_{p_{x}}$ and $\Delta_{p_{y}}$ with $\Delta=0.1$ eV.
Hall conductivity in the chiral phase–– The chiral $E^{\prime}$ state given by the $\Delta_{p+ip}$ pairing breaks time-reversal symmetry, allowing a finite optical Hall conductivity $\sigma_{xy}^{H}=-\sigma_{yx}^{H}$. Fig. 4 shows a comparison of the optical Hall and regular conductivities in the chiral $E^{\prime}$ state with the pairing $\Delta_{p+ip}$ in the frequency range of the superconducting gap. The Hall conductivity is comparable to the interband optical responses $\sigma_{xx}=\sigma_{yy}$ in the material, but notably while both $\sigma_{xx}$ and $\sigma^{H}_{xy}$ exhibit a peak at around $1.6\Delta$ associated with the maximal gap size and a large JDOS of the electron and hole pockets along the $K\Gamma$ high-symmetry line (see Fig. 2(a)), the size of this peak is much smaller for $\sigma^{H}_{xy}$. At a second peak around $3.6\Delta$, however, exhibits a similar magnitude in $\sigma_{xx}$ and $\sigma^{H}_{xy}$ associated with the higher-frequency van Hove singularity along the $K\Gamma$ high-symmetry line (see Fig. 2 (c), dotted-dashed line). As explained in Appendix II.1, the anomalously small peak in $\sigma_{xy}^{H}$ occurs due to a cancellation of matrix elements from transitions originating from different bands.
@@
Figure 4: Hall conductivity $\sigma^{H}_{xy}$ of the chiral superconducting state $\Delta_{p+ip}$ with $\Delta=0.1$ eV at low (main figure) and high (inset) frequencies. The diagonal component of the optical conductivity $\sigma_{xx}$ is shown for comparison at low frequencies.
Discussion–– In our work, we have predicted the optical conductivity of single layers of an H polytype of a TMD with $D_{3h}$ symmetry. These results most directly apply to single layers of NbSe2, where signatures of a nematic pairing channel have been observed in Refs. [Hamill et al., 2021] and [Cho et al., 2022]. Our results also bear relevance to 4Hb-TaS2, where superconductivity is dominated by the H layers in the T/H bulk structure. It should be noted this structure has an extra inversion center between the H layers and there are therefore two possibilities for inversion-even and odd versions of the $E^{\prime}$ state, which correspond to irreps $E_{2g}$ and $E_{1u}$ of the group $D_{6h}$ and should remain nearly degenerate. Finally, we speculate that our results may also be relevant to the superconducting state observed in CrBr3/NbSe2Kezilebieke et al. (2020), which coexists with ferromagnetism and shows chiral edge modes corresponding to a topological chiral superconductor.
[…]
In summary, we expect our work will help establish whether unconventional superconductivity in different TMD platforms is due to $E^{\prime}$ pairing, taking advantage of the multiorbital nature of their band structure with large Ising SOC which leads to unique optical conductivity signatures.
I Acknowledgements
F. J. is supported by Grant PID2021-128760NB0-I00 from the Spanish MCIN/AEI/10.13039/501100011033/FEDER, EU. M.-Á.S. was supported by EPSRC grant EP/X012239/1 and Programa Red Guipuzcoana de Ciencia Tecnología e Innovación No. 2021-CIEN-000070-01 Gipuzkoa Next.
II End Matter
II.1 Brillouin Zone distribution of the Hall response in $\Delta_{p+ip}$
In the chiral superconducting phase, the low-energy peak of the optical conductivity $\sigma_{xx}$ and the onset of the Hall conductivity is associated with the gap size, as discussed in the main text. The $\sigma_{xx}$ component exhibits a large peak around $1.6\Delta$ associated with the large JDOS between the closest bands to the Fermi level along the $K\Gamma$ high-symmetry lines and the positive-defined optical conductivity integrand over the Brillouin Zone (see Fig. 5 (a)). However, in the case of $\sigma^{H}_{xy}$ the velocity matrix elements connecting the closest bands to the Fermi level over the Brillouin Zone can be positive or negative (see Fig. 5 (b)). This leads to a cancellation of contributions that suppresses the peak that is present in the $\sigma_{xx}$ component due to the positive value of its velocity matrix elements over the whole Brillouin Zone.
The secondary peak at $\omega\sim 3.6\Delta$, associated with higher-frequency Van Hove singularity along $K\Gamma$ (see Fig. 2 (c), dashed-dotted line) is similar in size for $\sigma_{xx}$ and $\sigma^{H}_{xy}$ (see Fig. 4). This is due to the dominance of positive velocity matrix elements contributing to the Hall conductivity in that range of frequencies, much more similar to the $\sigma_{xx}$ case (Figs. 5 (c) and (d)) than in the primary peak case. The positive-integrand predominance in the secondary peak is due to the large positive contribution of the Van Hove singularity along $K\Gamma$, while having a limited presence of the negative-contributing bands (Fig. 5 (d)), which were more relevant in the primary peak energy range (Fig. 5 (b)).
@@ II.1 Brillouin Zone distribution of the Hall response in $\Delta_{p+ip}$
Figure 5: Brillouin Zone-resolved optical conductivity $\sigma_{xx}$ (left column) and $\sigma^{H}_{xy}$ (right) column integrated between $\omega=1.3\Delta$ and $\omega=1.8\Delta$ (upper row) corresponding to the low-energy peak in the response in Fig. 4, and $\omega=3\Delta$ and $\omega=4\Delta$ (lower row) corresponding to the secondary peak in the response in Fig. 4.

### v2

@@ IV Optical conductivity
${\rm Im}\sigma^{H}_{xy}=\frac{\pi e^{2}}{2\omega V}\sum_{m\neq n}\mathrm{Im}\left[v_{nm}^{x}v_{mn}^{y}\right]f_{nm}\delta\left(\epsilon_{n}-\epsilon_{m}+\omega\right),$ (9)
@@
where $e$ is the electron charge, $V$ is the volume of the system, $f_{nm}=f(\varepsilon_{n})-f(\varepsilon_{m})$ is the difference of Fermi functions, which depend on the energy $\varepsilon_{n}=E_{n}-\mu$ of band $n$, the chemical potential $\mu$ and the inverse temperature $\beta=1/k_{B}T$ where $k_{B}$ is the Boltzmann constant. We set $\beta=300\mathrm{eV}^{-1}$ ($T\simeq 3.87$K). The velocity matrix elements $v^{i}_{mn}$ account for the transition amplitude between states $n,m$ and are defined as:
@@
Note that the optical conductivity of clean multiband superconductorsAhn and Nagaosa (2021a) as computed here can be reduced in size for some types of superconductors—but not suppressed—by the inclusion of vertex correctionsWatanabe and Watanabe (2024), which are outside the scope of this work.
V Anisotropy in the nematic phase
@@
In the presence of $C_{3}$ symmetry, the optical conductivity satisfies $\sigma_{xx}=\sigma_{yy}$. Since the two-component nematic pairing breaks this symmetry, we consider the difference $\sigma_{xx}-\sigma_{yy}$ as a probe of the nematic pairing. Fig. 3 shows this difference in the range of frequencies near the superconducting gap for both $\Delta_{p_{x}}$ and $\Delta_{p_{y}}$. A sizable conductivity occurs due to transitions near the Fermi level, with significant differences between $\Delta_{p_{x}}$ and $\Delta_{p_{y}}$ that can help determine the pairing anisotropy.
@@
As a further optical signature of the pairing in multiorbital systems, we also consider higher frequency effects that occur due to transitions from the Fermi level BdG quasiparticles to unoccupied electron states, and from occupied electron states to BdG quasiparticles. These occur in the frequency range where the optical conductivity of the normal state is finite (see Fig. 1 (c)) and represent the transfer of spectral weight due to superconductivity as discussed in Ref. [Ahn and Nagaosa, 2021b]. However, here we consider responses which are zero in the absence of pairing, due to $C_{3}$ and time-reversal symmetry. The observation of any optical conductivity is therefore a signature of multiorbital pairing, which is shown in the inset of Fig. 3. The anisotropy of the nematic pairing is reflected in this spectrum as well.
@@
Figure 3: Anisotropy of the absorptive part of the optical conductivity at low (main figures) and high (inset) frequencies for the nematic phases $\Delta_{p_{x}}$ and $\Delta_{p_{y}}$ with $\Delta=0.1$ eV.
VI Hall conductivity in the chiral phase
The chiral $E^{\prime}$ state ($\Delta_{p+ip}$) breaks TRS, enabling a finite optical Hall response $\sigma_{xy}^{H}$. Figure 4 compares the absorptive components $\mathrm{Re}\sigma_{xx}$ and $\mathrm{Im}\sigma_{xy}^{H}$ across the gap window. The $\mathrm{Re}\sigma_{xx}$ component exhibits a large peak around $1.6\Delta$ associated with the large JDOS between the closest bands to the Fermi level along the $K\Gamma$ high-symmetry lines and the positive-defined optical conductivity integrand over the Brillouin Zone (see Fig. 5 (a)). However, in the case of $\mathrm{Im}\sigma^{H}_{xy}$ the velocity matrix elements connecting the closest bands to the Fermi level over the Brillouin Zone can be positive or negative (see Fig. 5 (b)). This leads to a cancellation of contributions that suppresses the peak that is present in the $\mathrm{Re}\sigma_{xx}$ component due to the positive value of its velocity matrix elements over the whole Brillouin Zone.
The secondary peak at $\omega\!\sim\!3.6\Delta$, associated with a higher‑frequency van Hove singularity along $K\Gamma$ [Fig. 2(c)], is similar in size for $\sigma_{xx}$ and $\sigma^{H}_{xy}$ [Fig. 4]. Here positive velocity‑matrix contributions dominate, resembling the longitudinal case [Figs. 5(c),(d)]. The predominance arises from the large positive weight of the van Hove feature along $K\Gamma$ and the reduced presence of negative contributions in this window [Fig. 5(d)], in contrast to the lower‑energy peak [Fig. 5(b)].
Figure 4: Hall response in the chiral state. We compare the anti-symmetric and symmetric absorptive components of the optical conducitivty, $\mathrm{Im}\sigma^{H}_{xy}$ and ${\rm Re}\sigma_{xx}$, respectively, for $\Delta_{p+ip}$ at $\Delta=0.1$ eV at low (main) and high (inset) frequencies. The shadowed region indicates the span of the gap across the Brillouin Zone.
@@ VI Hall conductivity in the chiral phase
Figure 5: Brillouin Zone-resolved absorptive part of the optical conductivity ${\rm Re}\sigma_{xx}$ (left column) and ${\rm Im}\sigma^{H}_{xy}$ (right) column integrated between $\omega=1.3\Delta$ and $\omega=1.8\Delta$ (upper row) corresponding to the low-energy peak in the response in Fig. 4, and $\omega=3\Delta$ and $\omega=4\Delta$ (lower row) corresponding to the secondary peak in the response in Fig. 4.
VII Experimental outlook
After establishing the general features of the band structure and optical conductivity for large values of $\Delta$ in the previous sections, we now focus on the measurability of the optical properties in the nematic and chiral phases. To do so, we will consider a realistic value of the superconducting order parameter $\Delta=1$ meV.
VII.1 Nematic phases
In the nematic phases $\Delta_{p_{x}}$ and $\Delta_{p_{y}}$, the small gap $\Delta=1$ meV results in two well-differentiated $\omega$ regions with different anisotropy, as we see in Fig. 6. The first region, spanning from $\omega=0$ eV to around $\omega=2$ eV features an anisotropy of the absorptive components of the optical conductivity $\mathrm{Re}(\sigma_{xx}-\sigma_{yy})\sim 10^{-7}e^{2}/\hbar$. The relevant quantity to consider experimental feasibility is the relative size of the anisotropy compared to the optical conductivity in that region, $\Delta\sigma/\sigma\equiv\frac{\mathrm{Re}(\sigma_{xx}-\sigma_{yy})}{\mathrm{Re}(\sigma_{xx}+\sigma_{yy})/2}\sim 10^{-6}$.
The second region (Fig. 6, inset) spans from $\omega=2.5$ eV to around $\omega=4.5$ eV. The anisotropy here is more than two orders of magnitude larger than that in the lower-energy region described in the previous paragraph, reaching values of $\mathrm{Re}(\sigma_{xx}-\sigma_{yy})\sim 5\times 10^{-5}e^{2}/\hbar$. The average value of the optical conductivity in this region is barely affected by the nematic phase, and its size is similar to that of the normal state in Fig. 1 (c). Therefore, the relative size of the anisotropy in this region is $\Delta\sigma/\sigma\sim 10^{-5}$.
The experimental feasibility of detecting optical anisotropies of magnitude $\Delta\sigma/\sigma\sim 10^{-6}$$10^{-5}$ is well-supported by current instrumentation standardsHarrison (2018); Thewalt et al. (2018). In the specific context of rotational symmetry breaking, reflectance anisotropy measurements have resolved small reflectivity features with a relative size of $\Delta R/R\sim 10^{-6}$ to characterize anisotropy in nematic superconductorsThewalt et al. (2018). Therefore, the expected anisotropy signature of the nematic superconducting phases discussed here for a realistic $\Delta=1$ meV is experimentally accessible with the current techniques.
Figure 6: Anisotropy of the absorptive components of the optical conductivity $\mathrm{Re}(\sigma_{xx}-\sigma_{yy})$ of the nematic superconducting states $\Delta_{p_{x}}$ and $\Delta_{p_{y}}$ for a realistic $\Delta=1$ meV in the $\omega\in[0,2.5]$ eV (main) and $\omega\in[2.5,4.5]$ eV (inset) regions.
VII.2 Chiral phase
The chiral $E^{\prime}$ phase produces a finite optical Hall conductivity $\sigma_{xy}^{H}$, which is most directly accessed via the polar Kerr effect (rotation of the reflected polarization) measured in a normal-incidence setup Argyres (1955); Kapitulnik et al. (2009). The Kerr angle is given byArgyres (1955); Kapitulnik et al. (2009)
$\theta_{\mathrm{K}}=-\mathcal{I}m\left[\frac{\tilde{n}_{\mathrm{L}}-\tilde{n}_{\mathrm{R}}}{\tilde{n}_{\mathrm{L}}\tilde{n}_{\mathrm{R}}-1}\right],$ (13)
where $\tilde{n}\equiv n+i\kappa$ is the complex refractive index for left $\tilde{n}_{\mathrm{L}}$ and right $\tilde{n}_{\mathrm{R}}$ circularly polarized light. In the case of a 2D sample with no substrate, the Kerr angle can be approximated byKapitulnik et al. (2009); Tse and MacDonald (2010)
$\theta_{K}\approx\tan\theta_{K}=\mathrm{Re}\left[\frac{\sigma_{xy}^{H}}{\sigma_{xx}}\right],$ (14)
where $\sigma_{xx}$ and $\sigma_{xy}^{H}$ are the complex conductivities, and we have used the fact that $\theta_{K}\ll 1$ in $\tan\theta_{K}\approx\theta_{K}$.
The Kerr angle for TaS2 with $\Delta=1$ meV in the chiral superconducting phase is shown in Fig. 7. Remarkably, the Kerr angle in this setup is in the $10^{-5}$$10^{-4}$ radians range, considerably larger than the resolution limit of $10^{-9}$ radians measured in other setups using Sagnac interferometryXia et al. (2006); Kapitulnik et al. (2009).
Figure 7: Kerr angle $\theta_{K}(\omega)$ for TaS2 with chiral $p+ip$ pairing with $\Delta=1$ meV. For $\omega<0.2$ eV the small size of $\Delta$ results in a large numerical noise and an absence of distinguishable features.
The large size of the Kerr rotation angle is due to the simultaneous scaling of both components, $\sigma_{xy}$ and $\sigma_{xx}$, with $\Delta$ in the described range. For values of $\Delta\in[1,10]$ meV we observe a strong peak at $\omega=330$ meV for all $\Delta$ values considered, indicating an interplay between superconductivity and the intrinsic features of the band structure that do not depend on the size of $\Delta$, leading to an observable feature in the Kerr angle. This is the same feature that gets displaced to $\omega=360$ meV for $\Delta=100$ meV, as discussed in Figs. 4 and 5, due to the large value of $\Delta$. This means that the cause of the persistent peak at $\omega=330$meV in Fig. 7 is the same physics discussed in Sec. VI.
Finally, we see that the size of the Kerr rotation angle in the Kerr-active region scales quadratically with $\Delta$ over a range of 1101-10 meV (see Fig. 8), providing an estimation of the order of magnitude of the angle for setups with similar properties and different superconducting pairings.
Figure 8: Persistent features and scaling of the Kerr angle for $\Delta\in[1,10]$meV. (a) Kerr angle $\theta_{K}(\omega)$ for a single H-layer of TaS2 with chiral $p+ip$ pairing for different values of $\Delta$. The position of the persistent peak at $\omega=330$ meV is indicated by the dashed gray line. For $\omega<0.2$ eV there is a strong numerical noise background covering any discernible features. (b) Scaling of the integrated Kerr angle $\theta_{K}\sim\Delta^{b}$ in the range $\omega\in[0,0.7]$ for different values of $\Delta$. The least squares fit of the shown data shows a quadratic exponent.
a
VIII Discussion
In our work, we have predicted the optical conductivity of single layers of an H polytype of a TMD with $D_{3h}$ symmetry. These results most directly apply to single layers of NbSe2, where signatures of a nematic pairing channel have been observed in Refs. [Hamill et al., 2021] and [Cho et al., 2022]. Our results also bear relevance to 4Hb-TaS2, where superconductivity is dominated by the H layers in the T/H bulk structure. It should be noted this structure has an extra inversion center between the H layers and there are therefore two possibilities for inversion-even and odd versions of the $E^{\prime}$ state, which correspond to irreps $E_{2g}$ and $E_{1u}$ of the group $D_{6h}$ and should remain nearly degenerate if the pairing is mainly intralayer. Finally, we speculate that our results may also be relevant to the superconducting state observed in CrBr3/NbSe2Kezilebieke et al. (2020), which coexists with ferromagnetism and shows chiral edge modes corresponding to a topological chiral superconductor.
Our quantitative outlook indicates that chiral $E^{\prime}$ order in H-TaS2 should produce Kerr rotations of $\theta_{K}\!\sim\!10^{-5}$ rad for realistic gap sizes, well within the sensitivity of modern Sagnac interferometryXia et al. (2006); Kapitulnik et al. (2009). Additionally, the nematic $E^{\prime}$ state yields a relative diagonal anisotropy $\mathrm{Re}(\sigma_{xx}-\sigma_{yy})/\sigma_{xx}\sim 10^{-5}$ without a Hall response, well within current experimental limits, enabling an optical symmetry diagnostic using polarization-resolved reflectivity. Likely, the experimental feasibility of the optical responses discussed here extends to other TMDs with superconducting gaps in the order of meV.
In summary, our results chart a concrete experimental pathway to help establish whether unconventional superconductivity in different TMD platforms is due to $E^{\prime}$ pairing: use linear optics to distinguish nematic versus chiral $E^{\prime}$ order, and quantify chiral TRSB via Kerr rotation.
IX Data and code availability
The code used to compute the optical conductivity and the generated datasets supporting this work will be made publicly available via Zenodo upon publication; the DOI will be inserted in the final version.
X Acknowledgements
F. J. is supported by Grant PID2021-128760NB0-I00 from the Spanish MCIN/AEI/10.13039/501100011033/FEDER, EU. M.-Á.S.-M. was supported by EPSRC grant EP/X012239/1 and Programa Red Guipuzcoana de Ciencia Tecnología e Innovación No. 2021-CIEN-000070-01 Gipuzkoa Next. D.M-S. acknowledges financial support from the Spanish Ministerio de Ciencia, Innovación y Universidades (MCIU) FPU fellowship No. FPU19/03195, and from the National Science Foundation (NSF) Materials Research Science and Engineering Centers (MRSEC) program through Columbia University under the Precision-Assembled Quantum Materials (PAQM) Grant No. DMR-2011738.
Appendix A Real and imaginary parts of the optical conductivity
The complete expression for the interband optical conductivity, containing real and imaginary parts, can be derived using standard Kubo formalism and reads
$\sigma_{ij}$ $=$ $\frac{\pi e^{2}}{2\omega V}\sum_{m\neq n}\frac{v^{i}_{nm}v^{j}_{mn}f_{nm}}{\varepsilon_{n}-\varepsilon_{m}+\omega+i\delta}$ (15)
$\underset{\delta\to 0}{\to}$ $\frac{\pi e^{2}}{2\omega V}\sum_{m\neq n}\left[\mathcal{P}-i\pi\delta(\varepsilon_{n}-\varepsilon_{m}+\omega)\right]$ (17)
$\times\left[\mathrm{Re}(v^{i}_{nm}v^{j}_{mn})+i\mathrm{Im}(v^{i}_{nm}v^{j}_{mn})\right]f_{nm},$
where we have used the Dirac identity
$\frac{1}{z+i\delta}\underset{\delta\to 0}{\to}\mathcal{P}\int_{-\infty}^{\infty}\mathrm{d}z\frac{1}{z}-i\pi\delta(z),$ (18)
where $\mathcal{P}$ is the Principal value of the integral and $\delta(z)$ is the Dirac delta function.
Taking the real part of Eq. (15) and noting that $v_{nm}^{i}v_{mn}^{i}=|v_{nm}^{i}|^{2}$ is a real quantity, we obtain the expression in Eq. (8) for the real part of the diagonal components of the optical conductivity.
Likewise, taking the imaginary part of Eq. (15) for the antisymmetric, off-diagonal terms $i\neq j$ we obtain the expression for the Hall conductivity in Eq. (9).
Appendix B Kramers-Kronig relations
The real and imaginary parts of the optical conductivity tensor $\mathrm{Re}\sigma+i\mathrm{Im}\sigma$ are related by the Kramers-Kronig identities:
$\mathrm{Re}\sigma(\omega)$ $=$ $\frac{1}{\pi}\mathcal{P}\int_{-\infty}^{\infty}\mathrm{d}x\frac{\mathrm{Im}\sigma(x)}{x-\omega},$ (19)
$\mathrm{Im}\sigma(\omega)$ $=$ $-\frac{1}{\pi}\mathcal{P}\int_{-\infty}^{\infty}\mathrm{d}x\frac{\mathrm{Re}\sigma(x)}{x-\omega}.$ (20)
Noting that $\mathrm{Re}\sigma(\omega)$ is even and $\mathrm{Im}\sigma(\omega)$ is odd in frequency, and following the standard derivation Dresselhaus , we obtain
$\mathrm{Re}\sigma(\omega)$ $=$ $\frac{2}{\pi}\int_{0}^{\infty}\mathrm{d}x\frac{x\mathrm{Im}\sigma(x)-\omega\mathrm{Im}\sigma(\omega)}{x^{2}-\omega^{2}},$ (21)
$\mathrm{Im}\sigma(\omega)$ $=$ $-\frac{2\omega}{\pi}\int_{0}^{\infty}\mathrm{d}x\frac{\mathrm{Re}\sigma(x)-\mathrm{Re}\sigma(\omega)}{x^{2}-\omega^{2}}.$ (22)
Since the bandwidth of the model used in the main text is finite, the absorptive components of $\sigma(\omega)$ computed in the main text are supported on a bounded frequency window and the integrals in Eq. (21) converge without ad-hoc truncations. To ensure convergence and numerical stability, the integrals in Eqs. (21) are computed for a maximum $\omega_{\mathrm{cutoff}}=50$eV.